# Derandomizing Multivariate Polynomial Factoring for Low Degree Factors

### v1

@@ 3.2 Factors of a constant degree product
3
4$(S_{\textup{Proj-Fac}},S_{\textup{Proj-Fac-mult}})=$ Projected-Factoring$(f,\delta)$
5for $(\widetilde{g},e)\in S_{\textup{Proj-Fac-mult}}$ do
@@
6
7      Compute $\widehat{g}=\Psi^{-1}(\widetilde{g})$, by Corollary 2.14 and its remark
8      Compute $g=\tau_{{\boldsymbol{\alpha}}}^{-1}(\widehat{g})$, and add $(g,e)$ to $L$
@@ 3.3 Constant degree factors
$\textup{T}_{{\rm Factor}(\mathcal{C}_{\mu}|_{{\mathcal{D}}_{\delta}})}=nd\,\textup{T}_{{\rm PIT}(\textup{Hom}[\mathcal{C}_{\mu}])}+d^{2}n^{{\rm poly}(\delta)}\,\textup{T}_{{\rm Div}(\partial\mathcal{C}_{\mu}/{\mathcal{D}}_{\delta})}+{\rm poly}(s,n^{{\rm poly}(\delta)},d).$
@@
Let $f({\boldsymbol{z}})\in\mathcal{C}$. To compute the factors of degree $\delta$ of $f$, we start again by invoking Projected-Factoring for $f$ to get $S$, the set of factors of the transformed trivariate polynomial $\Psi(\tau_{{\boldsymbol{\alpha}}}(f))$. As in the proof of Theorem 3.1, we compute the inverses of the transformations, $g=\tau_{{\boldsymbol{\alpha}}}^{-1}(\Psi^{-1}(\widetilde{g}))$.
@@
$L=\emptyset$, $L^{\prime}=\emptyset$      /* Initialize output list and intermediate candidates list */
3
$(S_{\textup{Proj-Fac}},S_{\textup{Proj-Fac-mult}})=$ Projected-Factoring$(f,\delta)$   /* Compute projected $3$-variate factors */
@@
5for $\widetilde{g}\in S_{\textup{Proj-Fac}}$ do
@@
6
7      Compute $\widehat{g}=\Psi^{-1}(\widetilde{g})$ (if the inverse exists) of degree $\leq\delta$ by Corollary 2.14
8      Compute $g=\tau_{{\boldsymbol{\alpha}}}^{-1}(\widehat{g})$
@@ 12      Let $z$ be a variable that $g$ depends on
13      Find the smallest $e\geq 1$ such that $g\not|\,\frac{\partial^{e}f}{\partial z^{e}}$ and add $(g,e)$ to list $L$
@@
As a consequence, we get a quasipolynomial-time algorithm to compute the constant-degree factors of a sparse polynomial. With a different proof, this was already shown by Kumar, Ramanathan, and Saptharishi [KRS24].
Corollary 3.3 (Constant-degree factors of sparse [KRS24]).
@@
$\textup{T}_{{\rm Factor}(\mathcal{C}_{\textup{sp}}|_{{\mathcal{D}}_{\delta}})}=(snd)^{{\rm poly}(\delta)\log s}.$
@@
Kumar, Ramanathan, and Saptharishi [KRS24] generalized sparse polynomials to polynomials computed by constant-depth circuits and showed that the constant-degree factors can be computed in subexponential time. We can now also derive this result via Theorem 3.2.
@@
$\mathcal{C}_{\textup{Depth-}{t}}(s,n,d)=\left\{\,p\in{\mathcal{P}}(n,d)\mid p\text{ has a circuit of size $s$ and depth }t\,\right\}.$
@@
    $\mathcal{C}_{\textup{Depth-}{t}}$ is closed for homogeneous components: For a polynomial $f\in\mathcal{C}_{\textup{Depth-}{t}}(s,n,d)$, all homogeneous components of $f$ are in $\mathcal{C}_{\textup{Depth-}{(t+1)}}(sd,n,d)$ (see [Oli16, Lemma 2.3]).
@@
    $\mathcal{C}_{\textup{Depth-}{t}}$ is closed under derivatives: For a polynomial $f\in\mathcal{C}_{\textup{Depth-}{t}}(s,n,d)$, a variable $z$ and $e\geq 1$, we have $\frac{\partial^{e}f}{\partial z^{e}}\in\mathcal{C}_{\textup{Depth-}{(t+1)}}(d^{2}\,s,n,d)$ (see [Oli16, Lemma 2.5]).
@@
Corollary 3.4 (Constant-degree factors of constant-depth [KRS24]).
@@
$\textup{T}_{{\rm Factor}(\mathcal{C}_{\textup{Depth-}{t}}|_{{\mathcal{D}}_{\delta}})}=\left(n\,(sd)^{O(t)}\right)^{O((sd)^{\epsilon})}\,.$
@@
$\textup{T}_{{\rm PIT}(\textup{Hom}[\mathcal{C}_{\textup{Depth-}{t}}])}=\textup{T}_{{\rm Div}(\partial\mathcal{C}_{\textup{Depth-}{t}}/{\mathcal{D}}_{\delta})}=\left(n\,(sd)^{O(t)}\right)^{O((sd)^{\epsilon})}\,.$
For the homogeneous components of $f$, this follows from property (ii) above that the components have again bounded-depth circuits. Hence, for PIT, we can apply Theorem 2.7.
For the divisibility test, this follows from property (i) and (iii) above and Corollary 2.9. ∎
@@ 3.4 Linear factors of $\rm ROABP$s
We can apply Theorem 3.2 in the case of read-once oblivious arithmetic branching programs, $\rm ROABP$s, and compute linear factors. $\rm ROABP$s are arithmetic branching programs (ABPs) where there is an order on the variables and on every path of the ABP, every variable is evaluated in this order and only once. As size measure for $\rm ROABP$s, usually the width $w$ is taken. Its size as a graph is then bounded by $nw$, where $n$ is the number of variables. Width and size can greatly vary depending on the variable order. We consider the any-order model. That is, when we say that a polynomial has an $\rm ROABP$ of width $w$, it is a bound for all orders of the variables. For a more detailed definition see for example [GKS17, GKST17]. Let
$\mathcal{C}_{\rm ROABP}(w,n,d)=\left\{\,p\in{\mathcal{P}}(n,d)\mid p\text{ hasan any-order $\rm ROABP$ of width }w\,\right\}.$
$\rm ROABP$s generalize sparse polynomials: $\mathcal{C}_{\textup{sp}}(s,n,d)\subseteq\mathcal{C}_{\rm ROABP}(s,n,d)$.
@@
Let $f\in\mathcal{C}_{\rm ROABP}(w_{1},n,d)$ and $g\in\mathcal{C}_{\rm ROABP}(w_{2},n,d)$. Then
@@
    $\textup{Hom}_{d}[f]\in\mathcal{C}_{\rm ROABP}((d+1)\,w_{1},n,d)$.
@@
    $\partial_{z^{e}}(f)\in\mathcal{C}_{\rm ROABP}(dw_{1},n,d-e)$, for a variable $z$ and $e\leq d$.
@@
    $(a_{0}+a_{1}z_{1}+a_{2}z_{2}+\cdots+a_{n}z_{n})^{d}\in\mathcal{C}_{\rm ROABP}(d+1,n,d)$, for $a_{0},a_{1},a_{2},\dots,a_{n}\in\mathbb{F}$.
@@
  • polynomial time ${\rm poly}(ndw)$, in the whitebox setting [RS05],
  • quasipolynomial time $(ndw)^{O(\log\log w)}$, in the blackbox setting [GKS17, GG20].
@@
Let $f({\boldsymbol{z}})\in\mathcal{C}_{\rm ROABP}(w,n,d)$ and $\ell({\boldsymbol{z}})$ be linear. We apply Lemma 2.8. Let $S=[2d^{2}+1]$ and find an ${\boldsymbol{\alpha}}\in\mathbb{F}^{n}$ such that $\ell({\boldsymbol{\alpha}})\neq 0$. Such an ${\boldsymbol{\alpha}}$ can be found in time $O(n)$. Let $c_{\beta,i}\in\mathbb{F}$ be constants such that $\ell|f\iff\widehat{f}=f({\boldsymbol{z}}+{\boldsymbol{\alpha}})-\ell({\boldsymbol{z}}+{\boldsymbol{\alpha}})\,\tilde{h}=0$, where
$\widetilde{h}=\sum_{\beta\in S}f(\beta{\boldsymbol{z}}+{\boldsymbol{\alpha}})\sum_{0\leq i\leq d}\,c_{\beta,i}\ell(\beta{\boldsymbol{z}}+{\boldsymbol{\alpha}})^{i}\,.$
We consider the width of an $\rm ROABP$ that computes $\widehat{f}$. The terms $\ell(\beta{\boldsymbol{z}}+{\boldsymbol{\alpha}})^{i}$ in $\widetilde{h}$ have $\rm ROABP$s of width $i+1\leq d+1$ by Lemma 3.5 (v). Applying Lemma 3.5 (i) and (ii), we get that $\widetilde{h}$ has an $\rm ROABP$ of width $O(wd^{4})$. Hence, for $\widehat{f}$, we get an $\rm ROABP$ of width $O(w^{2}d^{4})$. Now the claim follows from Lemma 3.6. ∎
@@
When the input $f$ is $s$-sparse, this result is already known due to Volkovich [Vol15, Theorem 4].
@@ 4 Computing the sparse factors
$\mathcal{C}_{\textup{sp}}(s,n,d)$ =$=\left\{\,p\in{\mathcal{P}}(n,d)\mid\textup{sp}(p)\leq s\,\right\}.$ (20)
@@
Given $f\in\mathcal{C}_{\textup{sp}}$, we can test whether $f$ is irreducible in time ${\rm poly}(snd)\,\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}$.
@@
The irreducibiltiy test first finds an ${\boldsymbol{\alpha}}\in(\mathbb{F}\setminus\{0\})^{n}$ such that $\textup{Hom}_{d}[f]({\boldsymbol{\alpha}})\neq 0$ in time ${\rm poly}(snd)$ by Theorem 2.6. Then, by solving ${\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})$, one can find a set $\mathcal{H}_{{\boldsymbol{\alpha}}}$ such that $f({\boldsymbol{z}})$ is reducible if and only if $f({\boldsymbol{\alpha}}x+{\boldsymbol{\beta}}t+{\boldsymbol{\gamma}})$ is reducible, for every $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in\mathcal{H}_{{\boldsymbol{\alpha}}}$. Whether a bivariate polynomial is reducible can be checked in time ${\rm poly}(d)$ by Lemma 2.4. ∎
@@
$\textup{T}_{{\rm Factor}(\mathcal{C}_{\mu}|_{\mathcal{C}_{\textup{sp}}})}=nd\,\textup{T}_{{\rm PIT}(\textup{Hom}[\mathcal{C}_{\mu}])}+{\rm poly}(snd)\,\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}\,\textup{T}_{{\rm Div}(\partial\mathcal{C}_{\mu}/\mathcal{C}_{\textup{sp}})}\,.$
@@
$\widehat{f}(x,t)=f({\boldsymbol{\alpha}}x+{\boldsymbol{\beta}}t+{\boldsymbol{\gamma}})$
@@
$\widehat{f}_{{\boldsymbol{\omega}}}(x,t_{1},t_{2})=f({\boldsymbol{\alpha}}x+{\boldsymbol{\beta}}t_{1}+({\boldsymbol{\omega}}-{\boldsymbol{\gamma}})t_{2}+{\boldsymbol{\gamma}})$
in dense representation by interpolation. Factor $g$ is mapped correspondingly to $\widehat{g}_{{\boldsymbol{\omega}}}(x,t_{1},t_{2})=g({\boldsymbol{\alpha}}x+{\boldsymbol{\beta}}t_{1}+({\boldsymbol{\omega}}-{\boldsymbol{\gamma}})t_{2}+{\boldsymbol{\gamma}})$. Note that $\widehat{f}_{{\boldsymbol{\omega}}}$ and $\widehat{g}_{{\boldsymbol{\omega}}}$ are monic polynomial in $x$ and $\widehat{g}_{{\boldsymbol{\omega}}}$ remains irreducible, since $\widehat{g}_{{\boldsymbol{\omega}}}(x,t,0)=g({\boldsymbol{\alpha}}x+{\boldsymbol{\beta}}t+{\boldsymbol{\gamma}})$ is irreducible, for a good point $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})$. We factorize $\widehat{f}_{{\boldsymbol{\omega}}}$ over $\mathbb{F}$ by Lemma 2.4. Let $F_{{\boldsymbol{\omega}}}=\{\widehat{g}_{{\boldsymbol{\omega}},1},\widehat{g}_{{\boldsymbol{\omega}},2},\ldots,\widehat{g}_{{\boldsymbol{\omega}},r}\}$ be the set of irreducible factors of $\widehat{f}_{{\boldsymbol{\omega}}}$. Then we have $\widehat{g}_{{\boldsymbol{\omega}}}\in F_{{\boldsymbol{\omega}}}$.
@@
$L=\emptyset$, $L^{\prime}=\emptyset$      /* Initialize output list and intermediate candidates list */
3
Find ${\boldsymbol{\alpha}}\in(\mathbb{F}\setminus\{0\})^{n}$ such that $\textup{Hom}_{d}[f]({\boldsymbol{\alpha}})\neq 0$      /* by Lemma 2.2 */
@@
5Compute $\mathcal{H}_{{\boldsymbol{\alpha}}}={\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})({\boldsymbol{\alpha}})$
6for each $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in\mathcal{H}_{{\boldsymbol{\alpha}}}$ do

### v2

@@ 3.2 Factors of a constant degree product
4
[…]
6for $(\widetilde{g},e)\in S_{\textup{Proj-Fac-mult}}$ do
@@
7
[…]
@@ 3.3 Constant degree factors
$\textup{T}_{{\rm Factor}(\mathcal{C}_{\mu}|_{{\mathcal{D}}_{\delta}})}=nd\,\textup{T}_{{\rm PIT}(\textup{Hom}[\mathcal{C}_{\mu}])}+d^{2}n^{{\rm poly}(\delta)}\,\textup{T}_{{\rm Div}(\partial\mathcal{C}_{\mu}/{\mathcal{D}}_{\delta})}+{\rm poly}(s,n^{{\rm poly}(\delta)},d).$ (35)
@@
Let $f({\boldsymbol{z}})\in\mathcal{C}$. To compute the factors of degree $\delta$ of $f$, we start again by invoking Projected-Factoring for $f$ to get $S$, the set of factors of the transformed trivariate polynomial $\Psi(\tau_{{\boldsymbol{\alpha}}}(f))$. As in the proof of Theorem 3.1, we compute the inverses of the transformations, $)g=\tau_{{\boldsymbol{\alpha}}}^{-1}(\Psi^{-1}(\widetilde{g}$)).
@@
31ex
$\emptyset, L=L^{\prime}=\emptyset$      /* Initialize output and intermediate candidates lists */
@@
[…]
6for $\widetilde{g}\in S_{\textup{Proj-Fac}}$ do
@@
7
[…]
@@ 13  Let $z$ be a variable that $g$ depends on
[…]
@@
As a consequence, we get a quasi-polynomial-time algorithm to compute the constant-degree factors of a sparse polynomial. With a different proof, this was already shown by Kumar, Ramanathan, and Saptharishi [kumar2024deterministic].
Corollary 3.3 (Constant-degree factors of sparse [kumar2024deterministic]).
@@
$\textup{T}_{{\rm Factor}(\mathcal{C}_{\textup{sp}}|_{{\mathcal{D}}_{\delta}})}=(snd)^{{\rm poly}(\delta)\log s}.$ (36)
@@
Kumar, Ramanathan, and Saptharishi [kumar2024deterministic] generalized sparse polynomials to polynomials computed by constant-depth circuits and showed that the constant-degree factors can be computed in subexponential time. We can now also derive this result via Theorem 3.2.
@@
[…]
@@
    $\mathcal{C}_{\textup{Depth-}{t}}$ is closed for homogeneous components: For a polynomial $f\in\mathcal{C}_{\textup{Depth-}{t}}(s,n,d)$, all homogeneous components of $f$ are in $\mathcal{C}_{\textup{Depth-}{(t+1)}}(sd,n,d)$ (see [oliveira2016factors, Lemma 2.3]).
@@
    $\mathcal{C}_{\textup{Depth-}{t}}$ is closed under derivatives: For a polynomial $f\in\mathcal{C}_{\textup{Depth-}{t}}(s,n,d)$, a variable $z$ and $e\geq 1$, we have $\frac{\partial^{e}f}{\partial z^{e}}\in\mathcal{C}_{\textup{Depth-}{(t+1)}}(d^{2}\,s,n,d)$ (see [oliveira2016factors, Lemma 2.5]).
@@
Corollary 3.4 (Constant-degree factors of constant-depth [kumar2024deterministic]).
@@
$\textup{T}_{{\rm Factor}(\mathcal{C}_{\textup{Depth-}{t}}|_{{\mathcal{D}}_{\delta}})}=\left(n\,(sd)^{O(t)}\right)^{O((sd)^{\epsilon})}\,.$ (38)
@@
$\textup{T}_{{\rm PIT}(\textup{Hom}[\mathcal{C}_{\textup{Depth-}{t}}])}=\textup{T}_{{\rm Div}(\partial\mathcal{C}_{\textup{Depth-}{t}}/{\mathcal{D}}_{\delta})}=\left(n\,(sd)^{O(t)}\right)^{O((sd)^{\epsilon})}\,.$ (39)
For the homogeneous components of $f$, this follows from property (ii) and Theorem 2.7.
For the divisibility test, this follows from property (i) and (iii) above, and Corollary 2.9. ∎
@@ 3.4 Linear factors of $\rm ROABP$s
We can apply Theorem 3.2 in the case of read-once oblivious arithmetic branching programs, $\rm ROABP$s, and compute linear factors. $\rm ROABP$s are arithmetic branching programs (ABPs) where there is an order on the variables and on every path of the ABP, every variable is evaluated in this order and only once. For $n$ variables ${\boldsymbol{z}}=(z_{1},z_{2},\dots,z_{n})$, we describe the order by a permutation $\sigma\in S_{n}$ on the indices, $(z_{\sigma(1)},z_{\sigma(2)},\dots,z_{\sigma(n)})$. As size measure for $\rm ROABP$s, usually the width $w$ is taken. Its size as a graph is then bounded by $nw$, where $n$ is the number of variables. Width and size can greatly vary depending on the variable order. Let
[…]
When we skip the variable order $\sigma$, it means that the width can be achieved by some order,
$\mathcal{C}_{\rm ROABP(w,n,d)}=\bigcup_{\sigma\in S_{n}}\mathcal{C}_{\rm ROABP_{\sigma}}(w,n,d)\,.$ (41)
$\rm ROABP$s generalize sparse polynomials: $\mathcal{C}_{\textup{sp}}(s,n,d)\subseteq\mathcal{C}_{\rm ROABP_{\sigma}}(s,n,d)$, for any $\sigma\in S_{n}$.
@@
Let $f\in\mathcal{C}_{\rm ROABP_{\sigma}}(w_{1},n,d)$ and $g\in\mathcal{C}_{\rm ROABP_{\sigma}}(w_{2},n,d)$. Then
@@
    $\textup{Hom}_{d}[f]\in\mathcal{C}_{\rm ROABP_{\sigma}}((d+1)\,w_{1},n,d)$,
@@
    $\partial_{z^{e}}(f)\in\mathcal{C}_{\rm ROABP_{\sigma}}(w_{1},n,d-e)$, for a variable $z$ and $e\leq d$,
@@
    $(a_{0}+a_{1}z_{1}+a_{2}z_{2}+\cdots+a_{n}z_{n})^{d}\in\mathcal{C}_{\rm ROABP_{\sigma}}(nd+d+1,n,d)$, for $a_{0},a_{1},a_{2},\dots,a_{n}\in\mathbb{F}$ and any $\sigma\in S_{n}$ [saxena2008diagonal].
@@
  • polynomial time ${\rm poly}(ndw)$, in the whitebox setting [raz2005deterministic],
  • quasi-polynomial time $(ndw)^{O(\log\log w)}$, in the blackbox setting [DBLP:journals/toc/GurjarKS17, guo2020improved].
@@
Let $f({\boldsymbol{z}})\in\mathcal{C}_{\rm ROABP_{\sigma}}(w,n,d)$ and $\ell({\boldsymbol{z}})$ be linear. We apply Lemma 2.8. Let $S=[2d^{2}+1]$ and find an ${\boldsymbol{\alpha}}\in\mathbb{F}^{n}$ such that $\ell({\boldsymbol{\alpha}})\neq 0$. Such an $\boldsymbol{\alpha}$} can be found in time $O(n)$. Let $c_{\beta,i}\in\mathbb{F}$ be constants such that $\ell|f\iff\widehat{f}=f({\boldsymbol{z}}+{\boldsymbol{\alpha}})-\ell({\boldsymbol{z}}+{\boldsymbol{\alpha}})\,\tilde{h}=0$, where
$\widetilde{h}=\sum_{\beta\in S}f(\beta{\boldsymbol{z}}+{\boldsymbol{\alpha}})\sum_{0\leq i\leq d}\,c_{\beta,i}\ell(\beta{\boldsymbol{z}}+{\boldsymbol{\alpha}})^{i}\,.$ (42)
We consider the width of an $\rm ROABP$ that computes $\widehat{f}$. The terms $\ell(\beta{\boldsymbol{z}}+{\boldsymbol{\alpha}})^{i}$ in (42) have $\rm ROABP$s of width $ni+i+1\leq nd+d+1$ by Lemma 3.5 (v) in any variable order, and hence, in particular in the order $\sigma$ of $f$. Applying Lemma 3.5 (i) and (ii), we get that $\widetilde{h}$ has an $\rm ROABP$ in order $\sigma$ of width $O(wnd^{4})$. Hence, for $\widehat{f}$, we get an $\rm ROABP$ in order $\sigma$ of width $O(w^{2}n^{2}d^{4})$. Now the claim follows from Lemma 3.6. ∎
@@
When the input $f$ is $s$-sparse, this result is already known due to Volkovich [volkovich2015deterministically, Theorem 4].
@@ 4 Computing the sparse factors
Given $f\in\mathcal{C}_{\textup{sp}}$, we can test whether $f$ is irreducible in time ${\rm poly}(snd)+\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}$.
@@
We first transform $f({\boldsymbol{z}})$ via Lemma 2.2 into a polynomial $f_{\boldsymbol{\alpha}}(x,{\boldsymbol{z}})$ that is monic in the new variable $x$. Note that ${\boldsymbol{\alpha}}\in(\mathbb{F}\setminus\{0\})^{n}$ can be computed in time ${\rm poly}(snd)$ by Theorem 2.6.
Then, by solving ${\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})$, one can find a set $\mathcal{H}_{{\boldsymbol{\alpha}}}$ such that $f({\boldsymbol{z}})$ is reducible if and only if $f({\boldsymbol{\alpha}}x+{\boldsymbol{\beta}}t+{\boldsymbol{\gamma}})$ is reducible, for every $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in\mathcal{H}_{{\boldsymbol{\alpha}}}$. Whether a bivariate polynomial is reducible can be checked in time ${\rm poly}(d)$ by Lemma 2.4. ∎
@@
$\textup{T}_{{\rm Factor}(\mathcal{C}_{\mu}|_{\mathcal{C}_{\textup{sp}}})}={\rm poly}(snd)\,(\textup{T}_{{\rm PIT}(\textup{Hom}[\mathcal{C}_{\mu}])}\,+\,\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}\,+\,\textup{T}_{{\rm Div}(\partial\mathcal{C}_{\mu}/\mathcal{C}_{\textup{sp}})})\,.$ (44)
@@
$\widehat{f}(x,t)=f({\boldsymbol{\alpha}}x+{\boldsymbol{\beta}}t+{\boldsymbol{\gamma}})$ (45)
@@
$\widehat{f}_{{\boldsymbol{\omega}}}(x,t_{1},t_{2})=f({\boldsymbol{\alpha}}x+{\boldsymbol{\beta}}t_{1}+({\boldsymbol{\omega}}-{\boldsymbol{\gamma}})t_{2}+{\boldsymbol{\gamma}})$ (46)
in dense representation by interpolation. Factor $g$ is mapped correspondingly to $\widehat{g}_{{\boldsymbol{\omega}}}(x,t_{1},t_{2})=g({\boldsymbol{\alpha}}x+{\boldsymbol{\beta}}t_{1}+({\boldsymbol{\omega}}-{\boldsymbol{\gamma}})t_{2}+{\boldsymbol{\gamma}})$. Note that $\widehat{f}_{{\boldsymbol{\omega}}}$ and $\widehat{g}_{{\boldsymbol{\omega}}}$ are monic polynomial in $x$ and $\widehat{g}_{{\boldsymbol{\omega}}}$ remains irreducible, since $\widehat{g}_{{\boldsymbol{\omega}}}(x,t,0)=g({\boldsymbol{\alpha}}x+{\boldsymbol{\beta}}t+{\boldsymbol{\gamma}})$ is irreducible, for a good point $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})$.
We factorize $\widehat{f}_{{\boldsymbol{\omega}}}$ over $\mathbb{F}$ by Lemma 2.4. Let $F_{{\boldsymbol{\omega}}}=\{\widehat{g}_{{\boldsymbol{\omega}},1},\widehat{g}_{{\boldsymbol{\omega}},2},\ldots,\widehat{g}_{{\boldsymbol{\omega}},r}\}$ be the set of irreducible factors of $\widehat{f}_{{\boldsymbol{\omega}}}$. Then we have $\widehat{g}_{{\boldsymbol{\omega}}}\in F_{{\boldsymbol{\omega}}}$.
@@
31ex
$\emptyset, L=L^{\prime}=\emptyset$      /* Initialize output list and intermediate candidates list */
@@
Find ${\boldsymbol{\alpha}}\in(\mathbb{F}\setminus\{0\})^{n}$ such that $f_{{\boldsymbol{\alpha}}}(x,{\boldsymbol{z}})$ is monic in $x$      /* by Lemma 2.2 */
5
6Compute $\mathcal{H}_{{\boldsymbol{\alpha}}}={\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})({\boldsymbol{\alpha}})$
7for each $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in\mathcal{H}_{{\boldsymbol{\alpha}}}$ do